# Proper scoring rules for estimation and forecast evaluation

### v1

@@ 3.1.2 Characterizations
The definition of weak continuity of the supergradient map is given in Appendix A. Theorem 19 need not be true when $\{\delta_{x}:x\in X\}\not\subseteq\mathcal{P}$ as exemplified by the quadratic score for Lebesgue densities on $\mathbb{R}^{d}$, see Section 3.3. It has divergence $d(\mathbb{P},\mathbb{Q})=\int(p(\mathbf{y})-q(\mathbf{y}))^{2}\,\mathrm{d}\mathbf{y}$ but is not a kernel score.
@@
Let $S:\mathcal{P}\times\mathbb{R}^{d}\to\overline{\mathbb{R}}$ be a translation invariant kernel score with the matrix $\mathbf{B}$ and the spectral measure $\mu$ as in Theorem 20. Then there exists a $\sigma$-finite measure $\rho$ on $(0,\infty)$ with $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$ and probability measures $\{\nu_{r}:r\in(0,1)\}$ on the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$ such that
@@
    If $S$ is homogeneous of degree $\alpha\in\mathbb{R}$, then $\alpha\in(0,2)$, $\mathbf{B}=\mathbf{0}$ and $\mu$ satisfies
@@ 3.1.3 Continuous ranked probability score (CRPS)
The CRPS is an intuitive, interpretable, and convenient measure of forecasting accuracy. It is distance-sensitive, that is, it assigns better scores to forecasts, which put more probability closer to the observations. In contrast, the logarithmic score assigns harsh penalties regardless of the overall closeness of the forecast to the observations if the forecast density at any of the observations is too low. The CRPS is essentially the $L^{2}$ distance between the CDFs of the forecast and the observations (see Example 11), and as the integral of the quadratic (Brier) score for binary outcomes, it elegantly connects to the conceptually simpler point forecasting setting. Indeed, it reduces to the absolute error loss when point forecasts are issued. Moreover, homogeneity implies that scaling the observations scales the predictive distribution accordingly even when the CRPS is away from the minimum. Furthermore, the CRPS exists for all forecasts with finite first moment, which means that the forecast measure need not necessarily have a density. It can be easily estimated by (14)using samples drawn from the forecast, or explicitly computed if the forecast distribution has finite support. This is especially useful in applications such as ensemble forecasting in meteorology, where forecasts are routinely issued as discrete predictive distributions consisting of the empirical distributions of the ensemble of predictions.
@@
Closed-form expressions for the CRPS have been obtained for many important parametric families such as the Gaussian and generalized extreme value (GEV) distributions (Taillardat et al., 2016). Closed-form expressions are also available for forecast CDFs which are linear splines (Gasthaus et al., 2019). In the absence of closed-form expressions, one can resort to estimating the score by sampling as in (14) or numerical integration techniques such as kernel quadrature (Adachi et al., 2025). Directly implementing (14) for the CRPS leads to an inefficient $\mathcal{O}(n^{2})$ time computation where $n$ is the number of samples drawn from the forecast. This can be improved to almost linear $\mathcal{O}(n\log n)$ time using an algebraically equivalent expression (Kelen et al., 2025):
@@
Pic et al. (2023) derived the minimax rate of convergence for distributional regression with the CRPS and show that it is achieved by using $k$-nearest neighbors ($k$-NN) or kernel smoothing. Surprisingly, this rate is of the same order as the minimax optimal rate for point prediction. Theoretical guarantees for estimation, model selection and convex aggregation have been obtained by Clement and Zaoui (2024).
@@ 3.2 Local scoring rules
Arguably, the most popular and important of all proper scoring rules is the logarithmic score given by $S_{\mathrm{log}}(\mathbb{P},\mathbf{y})=-\log p(\mathbf{y})$, compare Examples 5 and 9. It is a local scoring rule of order 00, and remarkably, the only differentiable scoring rule this kind (Bernardo, 1979). Empirical risk minimization for the logarithmic score is equivalent to maximum likelihood estimation which is the foundation of classical parametric inference. The unparalleled historical success of maximum likelihood is in great part because it furnishes closed-form estimates for many important parametric families of distributions and is asymptotically optimal under rather mild regularity conditions.
@@
Theorem 12 can be used to derive a simple proof that the logarithmic score is the only differentiable local scoring rule of order 00. Indeed, let $S(\mathbb{P},\mathbf{y})=s(\mathbf{y},p(\mathbf{y}))$ be such a scoring function. The function $s(\mathbf{y},z)$ is differentiable in $z$, and its entropy is $H(\mathbb{P})=\int s(\mathbf{x},p(\mathbf{x}))p(\mathbf{x})\,\mathrm{d}\mathbf{x}$. By Equations (7) and (9), we must have
@@
with $z_{1}=(\mathrm{d}/\mathrm{d}y)\log p(y)=p^{\prime}(y)/p(y)$ and $z_{2}=(\mathrm{d}^{2}/\mathrm{d}y^{2})\log p(y)=p^{\prime\prime}(y)/p(y)-(p^{\prime}(y)/p(y))^{2}$. It is strictly proper for distributions $\mathbb{P}$ on $\mathbb{R}$ with densities $p$ that satisfy some regularity conditions (Ehm and Gneiting, 2012, Definition 2.3). Parry (2016) has studies local scoring rules in the multivariate case, and introduces extensive scoring rules which are multivariate scoring rules which decompose into sums of univariate scoring rules for individual components and provides a characterization. Dawid et al. (2012) characterizes scoring rules for general discrete outcome spaces which are local in that they depend only on the values of the probability mass function on a neighborhood of $x$ in a given undirected graph.
@@ 3.3 $f$-Scores, $gf$-scores and generalized kernel scores
Such scoring rules were introduced by Grünwald and Dawid (2004) as separable Bregman scores, although we refer to them as $f$-scores to avoid confusion with Bregman loss functions and divergences. Pérez (1967) introduced the entropies at (15) as generalized $f$-entropies. However, they are different from the $f$-entropies of Csiszár (1972). $f$-scores are the only proper scoring rules of the form $S(\mathbb{P},y)=g(p(y))-l(\mathbb{P})$ (Dawid, 2007).
@@
Another twist of the idea is to replace the integral in (17) with a different entropy, for example, the entropy of a kernel score as at (11), that is, considering entropies of the form
@@ 3.4 Proper scoring rules induced by elicitable functionals
for all $z\in\mathcal{Z}$, $\mathbb{P}\in cP$, where $Y\sim\mathbb{P}$, subject to the existence of expectations, with equality if and only if $z=T(\mathbb{P})$. Such a scoring function $s$ is called strictly consistent for the functional $T$. Relevant functionals that are elicitable are the mean, quantiles, expectiles, and ratios of expections. Some important functionals like the variance are not elicitable. There is a large literature on elicitability with possible starting points being Lambert et al. (2008); Gneiting (2011); Fissler and Ziegel (2016).
@@ 5 Applications
In this section, we try to review the many ways in which proper scoring rules can be used to address computational and statistical considerations in estimation and forecast evaluation contexts.
@@ 5.1.1 Non-normalized Densities
Although the logarithmic score yields closed-form expressions for many standard parametric families, it is often intractable for more complicated ones. This is often the case in statistics (see for example de Fondeville and Davison, 2018) and especially in some modern machine learning applications such as generative diffusion models (Song and Ermon, 2019), where we need to model probability densities more flexibly, say as $p_{\theta}(\mathbf{x})\propto\exp\eta_{\theta}(\mathbf{x})$, where $\eta_{\theta}$ could be a neural network. Estimating $\theta$ via maximum likelihood for such complicated models would require evaluating the normalization constant $c_{\theta}=\int\exp\eta_{\theta}(\mathbf{x})\,\mathrm{d}\mathbf{x}$, which is prohibitively expensive even for moderately high dimensions. The Hyvärinnen score allows for a much more tractable estimation of such models by circumventing the need to evaluate the normalization constant $c_{\theta}$ and reducing the problem to minimizing the empirical risk corresponding to
@@
which depends only on $\eta_{\theta}$. In this way, it complements methods like Langevin dynamics, which allow sampling from $p_{\theta}$ using only $\eta_{\theta}$. This makes it particularly convenient for many statistical applications where normalization can be an issue such as estimation of high-dimensional graphical models (Lin et al., 2016), causal discovery (Rolland et al., 2022) and estimation of point processes (Cao et al., 2024). In Bayesian statistics, the Hyvärinen score has been has been used for likelihood-free inference (Pacchiardi and Dutta, 2022), model selection (Dawid and Musio, 2015), the derivation of objective priors (Leisen et al., 2020), model comparison (Shao et al., 2019) and sampling from Markov chain Monte Carlo and variational inference approaches (Zhang et al., 2018; Modi et al., 2023). The recent emergence of diffusion models as a powerful approach to generative modeling of images, audio, video and text, is also partly due to advancements in the implementation of score matching (see for example Dhariwal and Nichol, 2021).
@@ 6 Limitations and extensions
While proper scoring rules versatile and important tools for estimation and forecast evaluation. However, they also have limitations. Proper scoring rules can not assess tail properties of distributions. For example, under weak conditions on the scoring rule $S$, Brehmer and Strokorb (2019) demonstrate that for a given distribution $\mathbb{P}$ on $\mathbb{R}$, one can always find a distribution $\mathbb{Q}$ such that the divergence $d(\mathbb{Q},\mathbb{P})=S(\mathbb{Q},\mathbb{P})-S(\mathbb{P},\mathbb{P})$ is arbitrarily small but $\mathbb{P}$ and $\mathbb{Q}$ are not tail equivalent. This negative result has further been strengthened by Allen et al. (2024).
In survival analysis, the classical approach to inference based on maximum likelihood (i.e. the logarithmic score) is complicated by the fact that observations are censored because this often renders the likelihood intractable. Consequently, many approaches based on alternative scoring rules such as the survival-CRPS (Avati et al., 2020) and integrated Brier score (Graf et al., 1999) have been devised over the years. Unfortunately, these have been found to be improper by Rindt et al. (2022), implying that they are unfit for evaluation or estimation. Rindt et al. (2022) establish the propreity of the classical right-censored log-likelihood and propose a tractable approach to survival analysis based on it. Moreover, Yanagisawa (2023) proposed proper extensions of the Brier and ranked probability scores to survival analysis.
@@
7 Acknowledgements
The authors would like to thank Sam Allen, Ritabrata Dutta, Sebastian Engelke, Rafael Frongillo, Tilmann Gneiting, Pierre Jacob, Alexander Jordan, Andrea Nava, Romain Pic, and Xinwei Shen for inspiring and fruitful discussions and for useful pointers to the literature.
@@ 7 Acknowledgements
Appendix A Proofs
The appendix contains proofs of formal statements in the article.
A.1 Proofs for Section 2
@@
The measure $\mu$ is supported on $\mathbb{R}^{d}\setminus\{\mathbf{0}\}$, which we can identify with $(0,\infty)\times\mathbb{S}^{d-1}$ via the transformations $r=\|\mathbf{u}\|\in(0,\infty)$ and $\sigma=\mathbf{u}/\|\mathbf{u}\|\in\mathbb{S}^{d-1}$. By Rokhlin’s measure disintegration theorem (see Simmons, 2012, Theorem 6.4), we can write
@@
    00 0 $=(c^{2}-c^{\alpha})(\mathbf{y}-\mathbf{m}_{\mathbb{P}})^{\top}\mathbf{B}(\mathbf{y}-\mathbf{m}_{\mathbb{P}}),$
    00 0 $=c^{\alpha}\int_{\mathbb{R}^{d}}|e^{i\mathbf{u}^{\top}\mathbf{y}}-f_{\mathbb{P}}(\mathbf{u})|^{2}\,\mathrm{d}\mu(\mathbf{u})-\int_{\mathbb{R}^{d}}|e^{ic\mathbf{u}^{\top}\mathbf{y}}-f_{\mathbb{P}}(c\mathbf{u})|^{2}\,\mathrm{d}\mu(\mathbf{u}),$
@@
    where we can write $\rho((0,1))=\mu(\mathbb{S}^{d-1}\times(0,1))$. In order for $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$ to hold, we must have $\alpha\in(0,2)$. This implies $\mathbf{B}=\mathbf{0}$.

### v2

@@ 3.1.2 Characterizations
The definition of weak continuity of the supergradient map is given in the appendix. Theorem 19 need not be true when $\{\delta_{x}:x\in X\}\not\subseteq\mathcal{P}$ as exemplified by the quadratic score for Lebesgue densities on $\mathbb{R}^{d}$, see Section 3.3. It has divergence $d(\mathbb{P},\mathbb{Q})=\int(p(\mathbf{y})-q(\mathbf{y}))^{2}\,\mathrm{d}\mathbf{y}$ but is not a kernel score, see Remark 3 in the appendix.
@@
Let $S:\mathcal{P}\times\mathbb{R}^{d}\to\overline{\mathbb{R}}$ be a translation invariant kernel score with the matrix $\mathbf{B}$ and the spectral measure $\mu$ as in Theorem 20. Then there exists a $\sigma$-finite measure $\rho$ on $(0,\infty)$ with $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$ and probability measures $\{\nu_{r}:r\in(0,\infty)\}$ on the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$ such that
@@
    If $S$ is homogeneous of degree $\alpha\in\mathbb{R}$, then either (a) $\alpha=2$ and $\mu$ is zero everywhere, or (b) $\alpha\in(0,2)$, $\mathbf{B}=\mathbf{0}$ and $\mu$ satisfies
@@ 3.1.3 Continuous ranked probability score (CRPS)
The CRPS is an intuitive, interpretable, and convenient measure of forecasting accuracy. It is distance-sensitive, that is, it assigns better scores to forecasts, which put more probability closer to the observations. In contrast, the logarithmic score assigns harsh penalties regardless of the overall closeness of the forecast to the observations if the forecast density at any of the observations is too low. The CRPS is essentially the $L^{2}$ distance between the CDFs of the forecast and the observations (see Example 11), and as the integral of the quadratic (Brier) score for binary outcomes, it elegantly connects to the conceptually simpler point forecasting setting. Indeed, it reduces to the absolute error loss when point forecasts are issued. Moreover, homogeneity implies that scaling the observations scales the predictive distribution accordingly even when the CRPS is away from the minimum. Furthermore, the CRPS exists for all forecasts with finite first moment, which means that the forecast measure need not necessarily have a density. It can be easily estimated by Equation 14 using samples drawn from the forecast, or explicitly computed if the forecast distribution has finite support. This is especially useful in applications such as ensemble forecasting in meteorology, where forecasts are routinely issued as discrete predictive distributions consisting of the empirical distributions of the ensemble of predictions.
@@
Closed-form expressions for the CRPS have been obtained for many important parametric families such as the Gaussian and generalized extreme value (GEV) distributions (Taillardat et al., 2016). Closed-form expressions are also available for forecast CDFs which are linear splines (Gasthaus et al., 2019). In the absence of closed-form expressions, one can resort to estimating the score by sampling as in Equation 14 or numerical integration techniques such as kernel quadrature (Adachi et al., 2025). Directly implementing Equation 14 for the CRPS leads to an inefficient $\mathcal{O}(n^{2})$ time computation where $n$ is the number of samples drawn from the forecast. This can be improved to almost linear $\mathcal{O}(n\log n)$ time using an algebraically equivalent expression (Kelen et al., 2025):
@@
Pic et al. (2023) derived the minimax rate of convergence for distributional regression with the CRPS and showed that it is achieved by using $k$-nearest neighbors ($k$-NN) or kernel smoothing. Surprisingly, this rate is of the same order as the minimax optimal rate for point prediction. Theoretical guarantees for estimation, model selection and convex aggregation have been obtained by Clement and Zaoui (2024).
@@ 3.2 Local scoring rules
Arguably, the most popular and important of all proper scoring rules is the logarithmic score given by $S_{\mathrm{log}}(\mathbb{P},\mathbf{y})=-\log p(\mathbf{y})$, compare Examples 5 and 9. It is a local scoring rule of order 00, and remarkably, the only differentiable scoring rule of this kind (Bernardo, 1979). Empirical risk minimization for the logarithmic score is equivalent to maximum likelihood estimation which is the foundation of classical parametric inference. The unparalleled historical success of maximum likelihood is in great part because it furnishes closed-form estimates for many important parametric families of distributions and is asymptotically optimal under rather mild regularity conditions.
@@
Theorem 12 can be used to derive a simple proof that the logarithmic score is the only differentiable local scoring rule of order 00. Indeed, let $S(\mathbb{P},\mathbf{y})=s(\mathbf{y},p(\mathbf{y}))$ be such a scoring function. The function $s(\mathbf{y},z)$ is differentiable in $z$, and its entropy is $H(\mathbb{P})=\int s(\mathbf{x},p(\mathbf{x}))p(\mathbf{x})\,\mathrm{d}\mathbf{x}$. By Equation (7) and using Equation (9), we must have
@@
with $z_{1}=(\mathrm{d}/\mathrm{d}y)\log p(y)=p^{\prime}(y)/p(y)$ and $z_{2}=(\mathrm{d}^{2}/\mathrm{d}y^{2})\log p(y)=p^{\prime\prime}(y)/p(y)-(p^{\prime}(y)/p(y))^{2}$. It is strictly proper for distributions $\mathbb{P}$ on $\mathbb{R}$ with densities $p$ that satisfy some regularity conditions (Ehm and Gneiting, 2012, Definition 2.3). Parry (2016) studied local scoring rules in the multivariate setting. Specifically, he introduced and characterized the so-called extensive scoring rules, which are multivariate scoring rules that decompose into sums of univariate scoring rules for individual components. Dawid et al. (2012) characterized scoring rules for general discrete outcome spaces which are local in that they depend only on the values of the probability mass function on a neighborhood of $x$ in a given undirected graph.
@@ 3.3 $f$-Scores, $gf$-scores and generalized kernel scores
Such scoring rules were introduced by Grünwald and Dawid (2004) as separable Bregman scores, although we refer to them as $f$-scores to avoid confusion with Bregman loss functions and divergences. Pérez (1967) introduced the entropies in Equation 15 as generalized $f$-entropies. However, they are different from the $f$-entropies of Csiszár (1972). $f$-scores are the only proper scoring rules of the form $S(\mathbb{P},y)=g(p(y))-l(\mathbb{P})$ (Dawid, 2007).
@@
Another twist of the idea is to replace the integral in Equation 17 with a different entropy, for example, the entropy of a kernel score as at Equation 11, that is, considering entropies of the form
@@ 3.4 Proper scoring rules induced by elicitable functionals
for all $z\in\mathcal{Z}$, $\mathbb{P}\in\mathcal{P}$, where $Y\sim\mathbb{P}$, subject to the existence of expectations, with equality if and only if $z=T(\mathbb{P})$. Such a scoring function $s$ is called strictly consistent for the functional $T$. Relevant functionals that are elicitable are the mean, quantiles, expectiles, and ratios of expections. Some important functionals like the variance are not elicitable. There is a large literature on elicitability with possible starting points being Lambert et al. (2008); Gneiting (2011); Fissler and Ziegel (2016).
@@ 5 Applications
In this section, we briefly review the many ways in which proper scoring rules can be used to address computational and statistical considerations in estimation and forecast evaluation contexts.
@@ 5.1.1 Non-normalized Densities
Although the logarithmic score yields closed-form expressions for many standard parametric families, it is often intractable for more complicated ones. This occurs frequently in statistics (see for example de Fondeville and Davison, 2018) and especially in some modern machine learning applications such as generative diffusion models (Song and Ermon, 2019), where we need to model probability densities more flexibly, say as $p_{\theta}(\mathbf{x})\propto\exp\eta_{\theta}(\mathbf{x})$, where $\eta_{\theta}$ could be a neural network. Estimating $\theta$ via maximum likelihood for such complicated models would require evaluating the normalization constant $c_{\theta}=\int\exp\eta_{\theta}(\mathbf{x})\,\mathrm{d}\mathbf{x}$, which is prohibitively expensive even for moderately high dimensions. The Hyvärinnen score allows for a much more tractable estimation of such models by circumventing the need to evaluate the normalization constant $c_{\theta}$ and reducing the problem to minimizing the empirical risk corresponding to
@@
which depends only on $\eta_{\theta}$. In this way, it complements methods like Langevin dynamics, which allow sampling from $p_{\theta}$ using only $\eta_{\theta}$. This makes it particularly convenient for many statistical applications where normalization can be an issue such as estimation of high-dimensional graphical models (Lin et al., 2016), causal discovery (Rolland et al., 2022) and estimation of point processes (Cao et al., 2024). In Bayesian statistics, the Hyvärinen score has been used for likelihood-free inference (Pacchiardi and Dutta, 2022), model selection (Dawid and Musio, 2015), the derivation of objective priors (Leisen et al., 2020), model comparison (Shao et al., 2019) and sampling from Markov chain Monte Carlo and variational inference approaches (Zhang et al., 2018; Modi et al., 2023). The recent emergence of diffusion models as a powerful approach to generative modeling of images, audio, video and text, is also partly due to advancements in the implementation of score matching (see for example Dhariwal and Nichol, 2021).
@@ 6 Limitations and Extensions
Proper scoring rules are versatile and important tools for estimation and forecast evaluation. However, they also have limitations. Proper scoring rules cannot assess tail properties of distributions. For example, under weak conditions on the scoring rule $S$, Brehmer and Strokorb (2019) demonstrate that for a given distribution $\mathbb{P}$ on $\mathbb{R}$, one can always find a distribution $\mathbb{Q}$ such that the divergence $d(\mathbb{Q},\mathbb{P})=S(\mathbb{Q},\mathbb{P})-S(\mathbb{P},\mathbb{P})$ is arbitrarily small but $\mathbb{P}$ and $\mathbb{Q}$ are not tail equivalent. This negative result has further been strengthened by Allen et al. (2024).
In survival analysis, the classical approach to inference based on maximum likelihood (i.e. the logarithmic score) is complicated by the fact that observations are censored because this often renders the likelihood intractable. Consequently, many approaches based on alternative scoring rules such as the survival-CRPS (Avati et al., 2020) and integrated Brier score (Graf et al., 1999) have been devised over the years, see also Dawid and Musio (2014). Unfortunately, these have been found to be improper by Rindt et al. (2022), implying that they are unfit for evaluation or estimation. Rindt et al. (2022) establish the propriety of the classical right-censored log-likelihood and propose a tractable approach to survival analysis based on it. Moreover, Yanagisawa (2023) proposed proper extensions of the Brier and ranked probability scores to survival analysis.
@@
Acknowledgements
The authors would like to thank an anonymous reviewer, Sam Allen, Philip Dawid, Ritabrata Dutta, Sebastian Engelke, Rafael Frongillo, Tilmann Gneiting, Pierre Jacob, Alexander Jordan, Kristof Kraus, Andrea Nava, Romain Pic, and Xinwei Shen for comments, inspiring and fruitful discussions, and useful pointers to the literature.
@@ Acknowledgements
Appendix A Proofs for Section 2
@@
Remark 1 (Derivation of Hyvärinen score entropy).
Clearly,
$H(\mathbb{P})$ $=\int\left[\Delta_{\mathbf{y}}\log p(\mathbf{y})+\frac{1}{2}\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|^{2}\right]p(\mathbf{y})\,\mathrm{d}\mathbf{y},$
$=\int[\nabla_{\mathbf{y}}\cdot\nabla_{\mathbf{y}}]\log p(\mathbf{y})p(\mathbf{y})\,\mathrm{d}\mathbf{y}+\frac{1}{2}\int\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|^{2}p(\mathbf{y})\,\mathrm{d}\mathbf{y}.$
Using the divergence theorem,
$\int[\nabla_{\mathbf{y}}\cdot\nabla_{\mathbf{y}}]\log p(\mathbf{y})p(\mathbf{y})\,\mathrm{d}\mathbf{y}$ $=\int\nabla_{\mathbf{y}}\cdot[\nabla_{\mathbf{y}}\log p(\mathbf{y})p(\mathbf{y})]\,\mathrm{d}\mathbf{y}-\int[\nabla_{\mathbf{y}}\log p(\mathbf{y})]\cdot[\nabla_{\mathbf{y}}p(\mathbf{y})]\,\mathrm{d}\mathbf{y}$
$=\oint\nabla_{\mathbf{y}}\log p(\mathbf{y})p(\mathbf{y})\,\mathrm{d}S(\mathbf{y})-\int[\nabla_{\mathbf{y}}\log p(\mathbf{y})]\cdot[\nabla_{\mathbf{y}}\log p(\mathbf{y})]p(\mathbf{y})\,\mathrm{d}\mathbf{y}$
$=0-\int\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|^{2}p(\mathbf{y})\,\mathrm{d}\mathbf{y}$
where the surface integral is zero because $\|\nabla_{\mathbf{y}}\log p(\mathbf{y})\|\to 0$ as $\mathbf{y}\to 0$. The conclusion follows.
@@
Remark 3 (Why the quadratic score on $\mathbb{R}^{d}$ is not a kernel score.).
Assume to the contrary. Then there exists a kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ such that
$\int(p(\mathbf{y})-q(\mathbf{y}))^{2}\,\mathrm{d}\mathbf{y}=\int k(\mathbf{y},\mathbf{y}^{\prime})(p(\mathbf{y})-q(\mathbf{y}))(p(\mathbf{y}^{\prime})-q(\mathbf{y}^{\prime}))\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}$
for all $p,q$. Since these are quadratic forms in $p-q$, we can use the parallelogram law to derive
$\int(p_{1}(\mathbf{y})-q_{1}(\mathbf{y}))(p_{2}(\mathbf{y})-q_{2}(\mathbf{y}))\,\mathrm{d}\mathbf{y}=\int k(\mathbf{y},\mathbf{y}^{\prime})(p_{1}(\mathbf{y})-q_{1}(\mathbf{y}))(p_{2}(\mathbf{y}^{\prime})-q_{2}(\mathbf{y}^{\prime}))\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}$
for any densities $p_{1},p_{2},q_{1},q_{2}$ on $\mathbb{R}^{d}$. Let $A$ be a compact subset of $\mathbb{R}^{d}$ and $\{e_{j}\}_{j=1}^{\infty}$ be an orthonormal basis of the space of functions $f:A\to\mathbb{R}$ satisfying $\int_{A}f(\mathbf{y})\,\mathrm{d}\mathbf{y}=0$. By taking
$p_{1}(\mathbf{y})=\frac{\max(e_{i}(\mathbf{y}),0)}{2\int|e_{i}(\mathbf{y})|\,\mathrm{d}\mathbf{y}},\quad q_{1}(\mathbf{y})=\frac{\max(-e_{i}(\mathbf{y}),0)}{2\int|e_{i}(\mathbf{y})|\,\mathrm{d}\mathbf{y}},\quad p_{2}(\mathbf{y})=\frac{\max(e_{j}(\mathbf{y}),0)}{2\int|e_{j}(\mathbf{y})|\,\mathrm{d}\mathbf{y}},\quad q_{2}(\mathbf{y})=\frac{\max(-e_{j}(\mathbf{y}),0)}{2\int|e_{j}(\mathbf{y})|\,\mathrm{d}\mathbf{y}},$
we derive
$\delta_{ij}=\int_{A}e_{i}(\mathbf{y})e_{j}(\mathbf{y})\,\mathrm{d}\mathbf{y}=\int_{A\times A}k(\mathbf{y},\mathbf{y}^{\prime})e_{i}(\mathbf{y})e_{j}(\mathbf{y}^{\prime})\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}$
for every $1\leq i,j<\infty$. Since $\{e_{i}(\mathbf{y})e_{j}(\mathbf{y}^{\prime})\}_{i,j=1}^{\infty}$ form a basis for square-integrable functions on $A\times A$, we have
$\int_{A\times A}|k(\mathbf{y},\mathbf{y}^{\prime})|^{2}\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}=\lim_{N\to\infty}\sum_{i,j=1}^{N}\left[\int_{A\times A}k(\mathbf{y},\mathbf{y}^{\prime})e_{i}(\mathbf{y})e_{j}(\mathbf{y}^{\prime})\,\mathrm{d}\mathbf{y}\,\mathrm{d}\mathbf{y}^{\prime}\right]^{2}=\lim_{N\to\infty}N=\infty.$
By Theorem 16, $k$ can be chosen to be a negative-definite kernel of the form $k(\mathbf{y},\mathbf{y}^{\prime})=-\langle\psi_{\mathbf{y}},\psi_{\mathbf{y}^{\prime}}\rangle$ which means $|k(\mathbf{y},\mathbf{y}^{\prime})|^{2}\leq|k(\mathbf{y},\mathbf{y})||k(\mathbf{y}^{\prime},\mathbf{y}^{\prime})|$. It follows that the entropy of the uniform density $\mathbf{1}_{A}(\mathbf{y})/\int_{A}\,\mathrm{d}\mathbf{y}$ is given by
$\int_{A}\|\psi_{\mathbf{y}}\|^{2}\frac{\,\mathrm{d}\mathbf{y}}{\int_{A}\,\mathrm{d}\mathbf{y}}=\int_{A}|k(\mathbf{y},\mathbf{y})|\frac{\,\mathrm{d}\mathbf{y}}{\int_{A}\,\mathrm{d}\mathbf{y}}=\infty.$
This contradicts the fact that the entropy of the quadratic score is finite for the uniform density. The conclusion follows.
@@
The measure $\mu$ is supported on $\mathbb{R}^{d}\setminus\{\mathbf{0}\}$, which we can identify with $(0,\infty)\times\mathbb{S}^{d-1}$ via the transformations $r=\|\mathbf{u}\|\in(0,\infty)$ and $\sigma=\mathbf{u}/\|\mathbf{u}\|\in\mathbb{S}^{d-1}$. By Rohlin’s measure disintegration theorem (see Simmons, 2012, Theorem 6.4), we can write
@@
    00 0 $=(c^{2}-c^{\alpha})(\mathbf{y}-\mathbf{m}_{\mathbb{P}})^{\top}\mathbf{B}(\mathbf{y}-\mathbf{m}_{\mathbb{P}}),$ (27)
    00 0 $=c^{\alpha}\int_{\mathbb{R}^{d}}|e^{i\mathbf{u}^{\top}\mathbf{y}}-f_{\mathbb{P}}(\mathbf{u})|^{2}\,\mathrm{d}\mu(\mathbf{u})-\int_{\mathbb{R}^{d}}|e^{ic\mathbf{u}^{\top}\mathbf{y}}-f_{\mathbb{P}}(c\mathbf{u})|^{2}\,\mathrm{d}\mu(\mathbf{u}),$ (28)
@@
    where we can write $\rho((0,1))=\mu(\mathbb{S}^{d-1}\times(0,1))$. In order for a nonzero measure $\rho$ to satisfy $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$, we must have $\alpha\in(0,2)$.
    From (27) and (28), we thus have two possibilities: (a) $\alpha=2$ implying $\alpha\not\in(0,2)$ and thus, $\mu$ is zero everywhere and (b) $\alpha\neq 2$ implying $\mathbf{B}=\mathbf{0}$ and $\mu$ has the above form with $\alpha\in(0,2)$ (without ruling out the possibility that it is zero everywhere).